40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10729     Entry time: Fri Nov 21 02:22:25 2014
Author: ericq 
Type: Update 
Category: LSC 
Subject: More ALS PRFPMI exploration 

 Similar to what Jenne did the other night, I kept the PRFPMI arm DoFs locked on ALS, in hopes to check out the RF error signals. 

I was able to stably sit at nominally zero offset in both CARM and DARM, tens of minutes at a time, and the PRMI could reacquire without a fuss. Arm powers would rest between 10 and 20, intermittently exhibiting the "buzzing" behavior that Jenne mentioned when passing through resonance. 100pm CARM offset means arm powers of around 10, so since our ALS RMS is on this order, this seems ok. I saw TRX get as high as 212 counts, which is just about the same that I've simulated as the maximum power in our IFO. 

To get this stable, I turned off all boosts on MICH and PRCL except PRCL FM6, and added matrix elements of 0.25 for TRX and TRY in the trigger line for the PRMI DoFs. The logic for this is that if the arm powers are higher than 1, power recycling is happening, so we want to keep things above the trigger down value of 0.5, even if POP22 momentarily drops. 

I also played around a bit with DARM offsets. We know from experience that the ALS IR resonance finding is not super precise, and thus zero in the CARM FM is not zero CARM offset when on ALS. The same obviously holds for DARM, so I moved the DARM offset around, and could see the relative strengths of flashes change between the arms as expected.

I've written down some GPS times that I'm going to go back and look at, to try to back out some information about our error signals. 

Lastly, there may be something undesirable happening with the TRX QPD; during some buzzing, the signal would fluctuate into negative values and did not resemble the TRY signal as it nominally would. Perhaps the whitening filters are acting up...

ELOG V3.1.3-